1. A 10uC point charge is separated by the distance of 28cm from 18uC charge. Calculate the magnitude and electrical force between two charges.


2. The electrical force between two point charges with the magnitude of 300uC and 500uc is 24N. How far apart are two charges?


1
Expert's answer
2022-03-31T11:48:38-0400

1.

"F=k\frac{q_1q_2}{r^2}"

"F=9*10^9*\frac{10*10^{-6}*18*10^{-6}}{0.28^2}=21\:\rm N"

2.

"r=\sqrt{\frac{kq_1q_2}{F}}"

"r=\sqrt{\frac{9*10^9*300*10^{-6}*500*10^{-6}}{24}}=7.5\:\rm m"